OpenStudy (anonymous):

arrange in increasing order 4/9,3/8,1/3

OpenStudy (lukebluefive):

Use a calculator, or divide by hand, to find the following approximations: \[\frac{1}{3}\approx 0.33, \frac{3}{8}=3.75, \frac{4}{9}=\frac{2}{3}\approx0.66\]